Can you please help me start this problem?
A car traveled at 50 miles per hour for 150 miles and then traveled at 75 miles per hour for 150 miles. What was the car’s average speed?

Found 2 solutions by Alan3354, stanbon:
Answer by Alan3354(30993)


A car traveled at 50 miles per hour for 150 miles and then traveled at 75 miles per hour for 150 miles. What was the car’s average speed?
-------------------
d = r*t
t = d/r
-----------
Find the times for the 2 segments.
Add them.
Add the two distances.
-----
d = r*t
r = d/t

Answer by stanbon(57361)

A car traveled at 50 miles per hour for 150 miles and then traveled at 75 miles per hour for 150 miles. What was the car’s average speed?
------
1st leg data:
rate = 50 mph ; distance = 150 miles ; time = 150/50 = 3 hrs
------
2nd leg data
rate = 75 mph : distance = 150 miles ; time = 150/75 = 2 hrs
-------
Average time = (total distance)/(total time) = 300 miles/5 hrs = 60 mph
=========================
cheers,
Stan H.
==========================